OpenStudy (samara8954):

A pair of shoes usually sells for $70. If the shoes are 30% off, and sales tax is 6%, what is the total price of the shoes, including tax? $48.23 $51.94 $52.43 $55.65

OpenStudy (samara8954):

@MARC_

OpenStudy (samara8954):

help

OpenStudy (anonymous):

sales price = $70 less 30% = $ 21 ======================= price $49 tax (6%) $2.49 ------------------------- Total $51.49

OpenStudy (samara8954):

is that answer correct

OpenStudy (anonymous):

nope

OpenStudy (samara8954):

than what is the answer

OpenStudy (anonymous):

\[(70+(70\times \frac{ 6 }{ 100 })=?\] \[\frac{ 371 }{ 5 }\times \frac{ 100-30 }{ 100 }=?\]

OpenStudy (anonymous):

solve it @Samara8954

OpenStudy (anonymous):

#Sorry bout that

OpenStudy (anonymous):

371/5 = 74.2 & (70/100) = 0.7 = 74.2 {+ 0.7) =74.9